OpenStudy (anonymous):

3 times the complement of an angel is equal to one half of its supplement how large is the angel

OpenStudy (anonymous):

u mean angle?

OpenStudy (anonymous):

yes sorry.

OpenStudy (johnweldon1993):

Well...you know complementary angles add to 90 degrees....and supplementary add to 180 degrees So 3 times the complement of an angle....let the angle = x....so (90 - x) would be the compliment.....3 times that = 3(90 - x) is equal to half the supplement of that angle...again the angle is x...that is (180 - x) so half that is ((180 - x)/2) Now set them equal to each other... \[\large 3(90 - x) =\frac{180 - x}{2}\] Multiply both sides by 2 \[\large 6(90 - x) = 180 - x\] Distribute that 6... \[\large 540 - 6x = 180 - x\] Add x to both sides... \[\large 540 - 5x = 180\] Subtract 540 from both sides... \[\large -5x = -360\] Divide both sides by -5 \[\large x = \frac{-360}{-5} = \space ? \]

OpenStudy (anonymous):

so its 72 .. thanks a lot.

OpenStudy (johnweldon1993):

No problem!